OpenStudy (bacon):

Please Help!! How do i find y in... 18*(4y/3)+15y=1026?

OpenStudy (anonymous):

18 x (4y/3) +15y = 1026 (4y/3)+ 15y = 1026 / 18 (4y + 15y) = 57 / 3 19y = 19 y = 1

OpenStudy (anonymous):

rmjeans, math error in step 2

OpenStudy (anonymous):

(4y/3)+15y=57, 4y+15y=57*3, 19y=171, y=171/19